pieces of moulding 7 1/2 feet long, 2 3/4 feet long and 71/2 feet long to frame a doorway. If started with 20 feet of moulding how much is left over?
Answer by rfer(16322)

2.25ft
the easiest way to do this problem is use decimals. Add them up and subtract from 20.
or find common denominator, and then subtract from 20
7 2/4+2 3/4+7 2/4=16 7/4=17 3/4
20-17 3/4= 2 1/4 ft